# Chiral order emergence driven by quenched disorder

### v2

@@ Abstract
Quenched disorder is expected to destroy order, such as when a random field is applied in a 2-dimensional magnetic system. Even when order exists in the presence of disorder, it is usually only the survival of the order of the clean model. We present here a surprising phenomenon where an order, driven by quenched disorder, emerges, that has nothing in common with the order present in the clean model. It’s a new type of order by disorder that differs from the usual thermal or quantum one. The classical Heisenberg model on the $J_{1}-J_{3}$ kagome lattice is studied by parallel tempering Monte Carlo simulations, with magnetic site vacancies. After analyzing the effect of a few impurities on the ground state, favoring non-coplanar configurations, we show the emergence of a low-temperature chiral phase and the progressive destruction of the collinear $q=4$ Potts order, the only order present in the clean system.
In frustrated two-dimensional magnets with continuous spins, discrete symmetries can be broken and lead to finite temperature phase transitions: time-reversal or lattice symmetry breaking induce chirality, stripe orders or valence bond crystals[1, 2]. Quenched disorder is expected to destroy or weaken such emerging order. However, on the classical kagome spin model with first and third neighbor Heisenberg exchanges ($J_{1}J_{3}$), we show that site disorder not only breaks the discrete $K_{4}$ order of the pure system[3], but but also brings about the appearance of a completely new chiral order (see Fig. 1). This phenomenon could turn out to be quite general, and is the consequence of two pieces that had previously been discovered but not assembled: i) Disorder can be the origin of order, as in the well-known case of thermal and quantum fluctuations[4], ii) Holes in a spin lattice favor the least-collinear configurations[5, 6].
@@
Figure 1: Phase diagram of the $J_{1}J_{3}$ model for $J_{1}=-1$ and $J_{3}=1$. $p$ is the rate of magnetic vacancies, $T$ is the temperature. In the blue region the chirality $\chi\neq 0$. In the green region the Potts parameter $\sigma\neq 0$. The dots and crosses indicate phase transitions.
In most cases the broken symmetries in a clean system can be determined from the set of ground states. When they are not all related by a symmetry of the Hamiltonian, we speak from accidental degeneracies, as in the case of classical spin liquids[7]. At finite temperature, entropic effects favor the subset of ground states with the highest density of low-energy excitations. This is called the thermal order by disorder (ObD) phenomena, which also occurs in its quantum version at zero temperature for finite spin values due to quantum fluctuations.
The kagome antiferromagnet is the paradigmatic model of ObD, embodied by the $S=1/2$ compound Herbertsmithite[8], where both thermal and quantum fluctuations select coplanar states among its extensive number of ground states[9, 10, 11]. Among the many other kagome compounds (Kapellasite[12], Volborthite[13], Haydeite[14, 15], Sr-Vesignieite[16, 16]…), Ba-Vesignieite BaCu3V2O8(OH)2[17, 18, 19, 20, 21] can be described by a classical spin model which also presents ObD, where we now study the effect of site dilution.
This $J_{1}J_{3}$ model has ferromagnetic Heisenberg interactions $J_{1}$ between nearest neighbors and significant antiferromagnetic ones $J_{3}$ between third neighbors[22] (see Fig. 2), with a Curie-Weiss temperature of about $-77K$:
@@
where $\mathbf{S}_{i}$ are 3d unit vectors, $\eta_{i}$ is equal to $1$ for a magnetic site and 00 for a non-magnetic site (vacancy). The $\eta_{i}$ are statistically independent (uncorrelated disorder). $p=1-\sum_{i}\eta_{i}/N$ is the dilution parameter, with $N$ the total number of sites. Site dilution is an example of quenched disorder. We recall that a phase transition in spin model with quenched disorder usually falls into one of the following two categories, depending on the order parameter and on the type of quenched disorder[23]: (i) random field disorder, which arises when the disorder lifts the degeneracy between different ordered states, explicitly breaking the symmetry. As a consequence of the Imry-Ma argument [24], this prevents long-range order in dimension $d\leq 2$, (ii) random $T_{c}$ disorder, which preserves the broken-symmetry phase, but introduces spatial variations in the critical temperature. According to the Aizenman-Wehr argument[25], a phase transition can still occur if the pure model satisfies $\nu d<2$, where $\nu$ is the correlation length exponent $\xi\sim|T-T_{c}|^{-\nu}$. A representative example of these two categories is the $J_{1}-J_{2}$ Ising model on the square lattice, where site dilution affects the low-temperature phases (which depend on the values of $J_{1}$ and $J_{2}$) differently. While it preserves the ferromagnetic phase (acting as a random $T_{c}$ disorder), it destroys the stripe phase (acting as a random field disorder), by inducing a local breaking of the $C_{4}$ rotational lattice symmetry[26, 23, 27].
Figure 2: Left: $J_{1}$ and $J_{3}$ interactions considered in the $J_{1}J_{3}$ model. Right: collinear $K_{4}$ order, with its unit cell represented by a dashed line.
Theoretical studies of the pure $J_{1}J_{3}$ model have revealed a complex phase diagram[3, 28]. For $J_{1}=0$, the lattice decouples into three independent square sublattices, each exhibiting either ferromagnetic ($J_{3}<0$) or antiferromagnetic ($J_{3}>0$) order, leading in a 3-sublattice (for $J_{3}<0$) or 6-sublattice order (for $J_{3}>0$) (see Fig. 3). Beyond the expected $O(3)$ global spin rotational symmetry, an additional accidental degeneracy exists. When $J_{3}>0$, this accidental degeneracy is not lifted by a small $J_{1}$ (whatever its sign). The ground state manifold is called the 3subAF phase (see Fig. 3). By setting $J_{3}=1$, the range of $J_{1}$ values where this phase survives is quite extended[3]: $-1.24\simeq-\frac{1+\sqrt{5}}{4}<J_{1}<1$. We fix the value of $J_{3}=1$ from now on. However, the aforementioned ObD lifts this degeneracy at low temperature, stabilizing a discrete collinear $K_{4}$ (or equivalently $q=4$ Potts) order. The emergence of such a discrete order parameter in two dimensions via ObD has been observed in other models, such as the XY or Heisenberg $J_{1}-J_{2}$ square lattice[29, 30], where two columnar states are selected from the ground state manifold, leading to a $\mathbb{Z}_{2}$ order. A similar phenomenon was recently reported in the $J_{1}-J_{2}-J_{3}$ square lattice[31] with an analogous $K_{4}$ symmetry breaking.
Figure 3: Spin orientation in the 3subAF phase on the kagome lattice. Circles indicates the three sublattices $A$, $B$ and $C$, related by the $J_{3}$ interaction (red, blue and green links). On each of the sublattices, spins have two opposite orientations, forming three independently oriented Néel orders. When all spins are collinear, we get the order of Fig. 2.
Inevitably, perturbations - such as spin anisotropies - are present in realistic systems and can lift the ground state degeneracy, potentially favoring phases with alternative broken symmetries. However, the $K_{4}$ phase of the J1J3HK can persist at moderate temperatures if its entropy advantage compensates the energy cost. We focus here on a specific perturbation: quenched disorder, which can induce unexpected effects[32, 33, 5], that are fundamentally different from uniform perturbations of the Hamiltonian. In the J1J3HK model, it occurs that a discrete order parameter, chirality, emerges from site disorder, because it favors non-collinear configurations. The time-reversal symmetry is broken at low temperature, leading to a finite temperature phase transition. This reveals an unexpected result: a chiral ordered phase arising from quenched disorder, which is fundamentally different from the $K_{4}$ phase stabilized by entropic effects. Importantly, this phenomenon differs from a mere symmetry reduction that selectively eliminates certain ground states[27].
@@
We present the phenomena selecting non-collinear configurations and then the numerical results evidencing the transition. Each configuration in the 3subAF phase is fully characterized by 3 spin orientations $(\mathbf{S}_{A},\mathbf{S}_{B},\mathbf{S}_{C})$ on a reference triangle $ABC$ of the lattice, as it allows to propagate these orientations and their opposites on the whole lattice (Fig. 3). Global spin rotation are symmetries of the model, thus ground states related by such transformations are considered as equivalent (this continuous symmetry is only broken at $T=0$ according to the Mermin Wagner theorem). The equivalence class of a ground state is is thus specified by a couple $(\bm{\sigma},\chi)$ of a three dimensional real vector $\bm{\sigma}$ and a chirality $\chi=\pm 1$ defined on a triangle $ABC$ as:
@@
Among these classes, 4 have collinear spins ±$\pm\mathbf{S}_{A}=\pm\mathbf{S}_{B}=\pm\mathbf{S}_{C}$ resulting in 4 possible values for $\bf\sigma$ and $\chi=0$ (represented on Fig. 2 right) and 2 others have orthogonal spins pointing to the vertices of an octahedron ($\bf\sigma=0$ and $\chi=\pm 1$). The configurations with orthogonal spins are called octahedral. These two chiral spin states (among all other chiral ones) have the particularity to not break any lattice symmetry[34].
@@
At low temperature and for $p=0$, the 4 collinear classes of the 3subAF states are favored through thermal ObD[3]. It induces a discrete symmetry breaking related to the choice of one of the four possible classes (emergent $q=4$ Potts, or $K_{4}$ order parameter). The order parameter for this order is denoted $\Sigma$ and is defined in [3], while the chiral order parameter is $\chi$: the average of $\chi_{ABC}$ over the lattice. Finite-size analysis have shown that the critical exponents are compatible with the $4$ state Potts model universality class[3].
@@
What are now the effect of a non-zero (small) $p$ in the $J_{1}J_{3}$ And first of all, does the $K_{4}$ order survive such disorder ? A random field coupled to a discrete order parameter is expected to destroy it in $d\leq 2$. In our case, although the disorder at hand is not a random field but site disorder, it still generates an effective random field, because both the $K_{4}$ order and the site vacancies break the translational lattice symmetries[23]. To illustrate this, we evaluate the effect of a small number of vacancies on the collinear spin configuration of Fig. 2 in the approximation where the state does not reorder to minimize its energy. It happens that a hole on a hexagon of green sites has the same energetic effect as a hole on a triangle of blue sites. However, as soon as two neighboring sites are empty, the energy difference with respect to the fully occupied state depends on the hole position: $8J_{3}-J_{1}$ for two sites of different colors versus $8J_{3}+J_{1}$ for two identical-color sites. Site disorder thus affects the 4 translationally equivalent collinear orders differently, similar to what a random field does on otherwise equivalent magnetizations. Thus, the $K_{4}$ order is expected to be killed by an infinitesimal $p>0$. We will see that this seductive scenario is challenged by the numerical simulations.
@@
Figure 4: Ground state spin-spin correlation $\mathbf{S}_{i}\cdot\mathbf{S}_{0}$ between sites $i$ of the $A$ and $C$ sublattices and the central site 00 belonging to the $B$ sublattice. The exchanges of the $J_{1}$ and $J_{3}$ links emanating from site 00 tend to zero. Correlations between $B$ sites are not shown (they are strongly antiferromagnetic and near ±$\pm 1$.
@@
Whatever, the competition between the 4 collinear states becomes completely irrelevant at very low temperature, as the energy minimization near vacancies kills any spin collinearity and allows the emergence of a new discrete order parameter, related to the chirality sign. Unlike $K_{4}$, chirality does not break any lattice symmetry and does not suffer from any random field effect, fully allowing a phase transition in the presence of site disorder. To illustrate this, we begin by considering a single vacancy on a $B$-site of the kagome lattice (Fig. 3). For $J_{1}=0$, the presence or absence of this spin has strictly no effect on the ground state manifold, but switching on $J_{1}$ lifts the degeneracy. The $A$ and $C$ spins of the four first neighbors of the impurity experience an apparent magnetic field of length $J_{1}$, in the direction opposite to the missing spin[35]. In a Néel antiferromagnet, parallel and perpendicular magnetic susceptibilities differ: $\mathcal{S}_{\parallel}=0$ while $\mathcal{S}_{\perp}>0$: the spins tend to orient perpendicularly to the magnetic field and to minimize their energy by a canting that decreases when moving away from the impurity. As a consequence, the $A$ and $C$ Néel ordered sub-lattices orient quasi-perpendicularly to the $B$ spins, with angles depending on the distance to the impurity (see correlations on Fig. 4) and alternating signs, similarly to Friedel oscillations already evidenced in other continuous spin models with single vacancies[36]. This phenomena is similar to the selection of the least-collinear order[5, 6], also observed on the $J_{1}-J_{2}$ square model[27, 37, 38], explained by an effective positive biquadratic interaction $(\mathbf{S}_{i}\cdot\mathbf{S}_{j})^{2}$ between neighboring spins. However, with a unique impurity, the $A$ and $C$ quasi-Néel orders remain mutually independent and we still expect them to align collinearly at inifinitesimal temperatures, a partial survival of the initial thermal ObD. With one impurity, we have only two perpendicular spin directions, which is not enough to specify a chirality. On the $J_{1}-J_{2}$ square lattice, the only 2 intricated Néel order prevent the apparition of chirality[38]. Here, with 3 sublattices and with vacancies on all types of sites, the 3 Néel orders tend to orient mutually perpendicularly, allowing to determine a local spin triedra with definite chirality.
@@
The phase diagram $T-p$ is displayed in Fig. 1: the red dot at $p=0$ corresponds to the already known phase transition of the pure model, belonging to the universality class of the 4-state Potts model[3] (with critical exponents verifying $\alpha/\nu=1$, $\gamma/\nu=7/4$, and $21/\nu=3/$ 2), as clearly shown by finite-size scaling analysis (FSSA). We use the following relations for the maxima of the specific heat $C_{V}$ and of the (connected) susceptibility $S_{X}$ of the order parameter $X$ (with $X=\Sigma$ or $\chi$) to determine the nature of the phase transitions[41]:
@@
$\begin{split}C_{V}^{\rm max}(L)&=\frac{\overline{\langle\mathcal{H}^{2}\rangle}-\overline{\langle\mathcal{H}\rangle^{2}}}{NT^{2}}\simeq aL^{\alpha/\nu}+b,\\ \mathcal{S}_{X}^{\rm max}(L)&=\frac{\overline{\langle X^{2}\rangle}-\overline{\langle X\rangle^{2}}}{NT}\simeq cL^{\gamma/\nu}+d,\\ T_{c}^{Y}(L)&\simeq eL^{-1/\nu}+T_{c}(\infty),\end{split}$ (3)
@@
where brackets denote thermal averages and overlines are averages over disorder realizations. $T_{c}^{Y}$ is the temperature of the maximum of $Y$, where $Y$ is $C_{V}$, $\Sigma$ or $\chi$.
@@
Figure 5: (left) Maxima of heat capacity $C_{V}^{\rm max}$ versus the linear lattice size $L$ and (right) $T_{c}^{C_{V}}(L)$ versus $L^{-3/2}$, for several dilution rates $p$.
@@
Figure 6: (left) Order parameter $\Sigma$ (increasing and decreasing with $T$) and $\chi$ (immediately decreasing) versus $T$ for different dilutions and $L=96$. (right) Maximum of the susceptibilty $S_{\Sigma}^{\rm max}$ versus $L^{1.75}$. The dotted lines correspond to linear fits.
We now discuss the phase diagram for $p\neq 0$. For low impurity rate $p<0.05$, the transition temperature between the paramagnetic phase and the $K_{4}$ order decreases rapidly with increasing $p$. We observe a linear dependence of $C_{V}^{\rm max}(L)$ (Fig. 5 left), but the amplitude $a$ decreases rapidly with $p$. The evolution of $T_{c}^{C_{V}}(L)$ (Fig. 5 right) confirms that $\nu=2/3$. Lastly, the maximum of the $K_{4}$ susceptibilty $\mathcal{S}_{\Sigma}^{\rm max}$ increases as $L^{7/4}$ with a slope decreasing with $p$ (Fig. 6 right). Thus FSSA clearly indicates that there is a phase transition in the same universality class as the pure model, i.e as the 4-state Potts model. This result seems to invalidate the argument that site dilution leads to the existence of local random fields, but we want to draw attention on the fact that other studies report transitions in presence of random fields, arguing that their lattice-symmetry breaking effect is very small[26].
@@
At lower temperatures, a second phase transition appears, where we observe simultaneously a rapid decrease of $\Sigma$ ($K_{4}$ order) and a rapid increase of $\chi$ (chiral order) as $T$ decreases (see Fig. 6 left). In the presence of site disorder, non-collinear configurations are energetically favored, overcoming the entropic advantage of the $K_{4}$ order at very low temperatures. The thermodynamic signature of this transition is the appearance of a small secondary maximum in $C_{V}$ at $T_{c}\simeq 4p$ for $p<0.05$, which increases weakly with $L$ (e.g., $C_{V}^{\rm max}(28)=1.15$, $C_{V}^{\rm max}(128)=1.20$). This prevents the determination of a reliable scaling exponent with FSSA, that however remains compatible with the scaling of the dilute Ising spin model[42]. $S_{\chi}$ appears to be weakly dependent on the system size $L$, while $S_{\Sigma}$ follows a scaling behavior with an exponent slightly larger than 1. To obtain stable results in this regime, a disorder average over $20$ realizations is required. Results are also also limited to $L\leq 80$, due to thermalization difficulties.

### v3

@@ Abstract
Quenched disorder can destroy magnetic order, for example when a random field is applied in a 2-dimensional Ising model. Even when an order exists in the presence of quenched disorder, it is usually only the survival of the order of the clean model. We present here a surprising phenomenon where an order emerges, driven by quenched disorder. This order has nothing in common with the order present in the clean model. This type of order by disorder differs from the usual thermal or quantum one. The classical $J_{1}-J_{3}$ Heisenberg model on the kagome lattice is studied by parallel tempering Monte Carlo simulations, with site dilution. After analyzing the effect of a few vacancies on the ground state, favoring non-coplanar configurations, we show the emergence of a low-temperature chiral phase and the progressive destruction of the collinear $q=4$ Potts order, the only order present in the absence of vacancies.
In two-dimensional frustrated magnets with continuous spin degrees of freedom, non-collinear or non-coplanar spin orders may break discrete symmetries, such as time-reversal or lattice symmetries. Then the Mermin-Wagner theorem allows for finite-temperature phase transitions, resulting in chiral orders, stripe patterns, or valence bond crystals [1, 2]. The symmetries of the low-temperature phase can usually be inferred from the structure of the ground-state manifold. When these states are not all related by the Hamiltonian’s symmetries, one speaks of accidental degeneracy, as observed in classical spin liquids [3]. At finite temperatures, entropic contributions favor ground states with a high density of low-energy excitations, a mechanism known as order by thermal disorder (ObTD)[4]. Similarly, quantum fluctuations (finite spin value) give rise to order by quantum disorder (ObQD). The selected states are typically the most-collinear spin configurations[5, 6, 7]. Several paradigmatic models illustrate the thermal or quantum ObD, among them the $J_{1}-J_{2}$ XYXY[6] or Heisenberg[8, 9] model on the square lattice where fluctuations favor collinear spin configurations, generating an Ising-like phase transition at finite temperature, the kagome antiferromagnet, where coplanar spin configurations are selected[10, 11, 12] and the $XY$ pyrochlore lattice[13], where a complex non-coplanar order is selected.
Quenched (structural) disorder, such as bond or site disorder, is generally expected to suppress or destabilize any kind of order. Phase transitions in spin models with quenched disorder typically fall into one of these two classes, according to the symmetry of the order parameter and the nature of the disorder [14]: (i) Random-field disorder, which explicitly breaks the symmetry between ordered states, typically precluding long-range order in dimensions $d\leq 2$ for a discrete order, $d\leq 4$ for a continuous order, by the Imry-Ma argument [15]; and (ii) Random-$T_{c}$ disorder, which preserves the symmetry of the ordered phase but induces spatial fluctuations in the critical temperature[16]. For example, in the 2d Ising model, random-fields suppress the transition[15], while site dilution (random-$T_{c}$ disorder) just reduces it, lowering the critical temperature[17, 18]. On the $J_{1}$$J_{2}$ Ising model on the square lattice, site dilution affects the low-temperature phases in distinct ways according to the order parameter, preserving ferromagnetic order (random-$T_{c}$ disorder) while destroying the stripe phase (random-field disorder), by locally breaking the $C_{4}$ lattice symmetry [19, 14, 20] and creating effective random fields.
However, when the order emerges from subtle phenomena such as ObTD or ObQD, quenched disorder not only destabilize it, but may induce the emergence of new orders. By opposition to ObTD or ObQD, quenched disorder is known to favor the least-collinear spin configurations[5, 6, 21, 7, 22, 23, 24, 25].
@@
Figure 1: Phase diagram of the $J_{1}J_{3}$HK model with $J_{1}=-1$ and $J_{3}=1$. $p$ denotes the dilution rate and $T$ the temperature. The blue region indicates a phase with non-zero scalar chirality $\chi\neq 0$, while the green region corresponds to non-zero $K_{4}$ order parameter $\sigma\neq 0$. Blue and red dots indicate phase transitions, while green crosses correspond to a crossover.
In this article, we present an $O(3)$ symmetric model where quenched disorder induces the emergence of a chiral phase: the $J_{1}-J_{3}$ Heisenberg model on the kagome lattice ($J_{1}J_{3}$HK model). This phase, spontaneously breaking the time-reversal symmetry at low temperatures, gives rise to a finite-temperature phase transition as soon as dilution is present, see Fig. 1. Remarkably, this phenomenon appears to be robust and potentially general in several Heisenberg models. This discovery is particularly noteworthy, as the resulting chiral phase is fundamentally distinct from the discrete 4-state Potts order (denoted $K_{4}$ in the following, as the Klein four-group) stabilized via ObTD[26]. Crucially, it does not merely arise from a symmetry reduction that selectively favors a subset of ground states [20], but represents a qualitatively different form of order induced by the presence of quenched disorder. Although, due to vacancies, the apparition of a chirality has already been reported in a $XY$ model in the seminal work of Henley[6], as well as the apparition of new phases in Heisenberg models with a magnetic field[7], this is to our knowledge the first mention of the emergence of a discrete chirality in a model with O(3) spin-rotation symmetry.
The $J_{1}J_{3}$HK model features ferromagnetic nearest-neighbor Heisenberg interactions ($J_{1}$) and antiferromagnetic third-neighbor couplings ($J_{3}$), as illustrated in Fig. 2. This model with no $J_{2}$ coupling yet large $J_{3}$ may appear unphysical and of purely theoretical interest. In fact, this model is the classical approximation of the $S=1/2$ $J_{1}J_{3}$HK model describing the Ba-Vesignieite compound BaCu3V2O8(OH)2[27, 28, 29, 30, 31] where the $J_{3}$ coupling dominates[32].
In the presence of site dilution, the Hamiltonian reads:
@@
where $\mathbf{S}_{i}$ are classical 3-component unit vectors, and $\eta_{i}=1$ or 0 indicates the presence or absence of a magnetic spin at site $i$, respectively. The $\eta_{i}$ are uncorrelated random variables representing quenched disorder. The dilution rate is defined as $p=1-\sum_{i}\eta_{i}/N$, where $N$ is the total number of sites.
Figure 2: Left: First ($J_{1}$) and third-neighbor ($J_{3}$) interactions in the $J_{1}J_{3}$HK model. Right: spin orientations of the collinear $K_{4}$ order, characterized by a four-site unit cell (dashed outline). Sites with the same spin orientation form hexagons (light green) and triangle (light blue).
@@
The $J_{1}J_{3}$HK model without site dilution exhibits a rich zero-temperature phase diagram [26, 33]. For $J_{1}=0$, the kagome lattice decouples into three independent (squeezed) and equivalent square sublattices (in different colors on the left of Fig. 3), each displaying ferromagnetic ($J_{3}<0$) or antiferromagnetic ($J_{3}>0$) ordering. This results in either a 3-sublattice or 6-sublattice configuration, depending on the sign of $J_{3}$ (see right of Fig. 3). Beyond global $O(3)$ spin rotation symmetry, the model hosts an accidental degeneracy which is not lifted by small $J_{1}$ values for $J_{3}>0$. Then the system resides in the so-called 3subAF phase. Fixing $J_{3}=1$, this phase remains stable over a wide interval: $1.24\lesssim J_{1}<1$ [26]. Each configuration in the 3subAF phase is fully characterized by 3 spin orientations $(\mathbf{S}_{A},\mathbf{S}_{B},\mathbf{S}_{C})$ on a reference triangle $ABC$ of the lattice, as it allows to propagate these orientations and their opposites on the whole lattice (Fig. 3). Global spin rotation are symmetries of the model, thus ground states related by such transformations are considered as equivalent (this continuous symmetry is only broken at $T=0$ according to the Mermin Wagner theorem). The equivalence class of a ground state is is thus specified by a couple $(\bm{\sigma},\chi)$ of a three dimensional real vector $\bm{\sigma}$ and a chirality $\chi=\pm 1$ defined on a triangle $ABC$ as:
@@
Among these classes, 4 have collinear spins $\pm\mathbf{S}_{A}=\pm\mathbf{S}_{B}=\pm\mathbf{S}_{C}$ resulting in 4 possible values for $\bf\sigma$ and $\chi=0$ (represented on Fig. 2 right) and 2 others have orthogonal spins pointing to the vertices of an octahedron ($\bf\sigma=0$ and $\chi=\pm 1$). The configurations with orthogonal spins are called octahedral. These two chiral spin states (among all other chiral ones) have the particularity to not break any lattice symmetry[34].
Figure 3: Spin structure in the 3subAF phase on the kagome lattice. Circles denote the three sublattices AA, $B$, and $C$, connected via third-neighbor $J_{3}$ couplings (colored links). Each sublattice independently exhibits Néel order with two opposing spin orientations (light/dark arrows, of the color of the sublattice). The collinear limit of this structure corresponds to the $K_{4}$ order shown in Fig. 2.
@@
At low temperature and for $p=0$, the 4 collinear classes of the 3subAF states are favored through ObTD[26]. It induces a discrete symmetry breaking related to the choice of one of the four possible classes (emergent $q=4$ Potts, or $K_{4}$ order parameter). The order parameter for this order is denoted $\Sigma$ and is defined from $\sigma$ (Eq. 2) similarly to the alternated magnetization of a Néel order[26], while the chiral order parameter is $\chi$: the absolute value of the average of $\chi_{ABC}$ over the lattice. Finite-size analysis have shown that the critical exponents are compatible with the $4$ state Potts model universality class[26]. Notably, there is a strong similitude with the $J_{1}-J_{2}$ Heisenberg model on the square lattice ($J_{1}J_{2}$HS), where the large-$J_{2}$ limit consists in 2 decoupled square lattices and only 2 collinear orders (distinct up to global spin rotations), leading to an Ising-like order.
@@
What are now the effect of a non-zero (small) $p$ in the $J_{1}J_{3}$HK model ? And first of all, does the $K_{4}$ order survive such quenched disorder ? First, we discuss the emergence of random fields destroying the $K_{4}$ order, then the apparition of chiral configurations and finally, we present the numerical simulations evidencing the chiral transition and the evolution of the $K_{4}$ phase transition. In the $J_{1}J_{3}$HK model, although the disorder at hand is not a random field per se but site disorder, it still generates an effective random field, because both the $K_{4}$ order and the site vacancies break the translational lattice symmetries[14]. To illustrate this, we evaluate the effect of a small number of vacancies on the collinear spin configuration of Fig. 2 in the approximation where the state does not reorder to minimize its energy. It happens that a hole on a hexagon of green spins has the same energetic effect as a hole on a triangle of blue spins. However, as soon as two neighboring sites are empty, the energy difference with respect to the fully occupied state depends on the hole position: $8J_{3}-J_{1}$ for two sites of different colors versus $8J_{3}+J_{1}$ for two identical-color sites. Site disorder thus affects the 4 translationally equivalent collinear orders differently, similar to what a random field does on otherwise equivalent magnetizations. According to the Imry-Ma argument, the $K_{4}$ order is thus expected to be killed by an infinitesimal $p>0$, as we are at the critical lower dimension $d=2$ for a discrete order[15].
@@
Figure 4: Ground state spin-spin correlation $\mathbf{S}_{i}\cdot\mathbf{S}_{0}$ between sites of the $A$ and $C$ sublattices and the central site 0 belonging to the $B$ sublattice, for a periodic $24\times 24\times 3$ lattice. The exchanges of the $J_{1}$ and $J_{3}$ links emanating from site 0 tend to zero. Correlations between $B$ sites are not shown (they are strongly antiferromagnetic and near $\pm 1$), and the diameter of circles saturate for absolute values above 0.04
@@
At very low temperature, the effective fields not only kill the $K_{4}$ order, but allows the emergence of a new discrete order parameter, related to the chirality sign. Unlike $K_{4}$, chirality does not break any lattice symmetry and does not suffer from any random field effect, fully allowing a phase transition in the presence of site disorder. To illustrate this, we begin by considering a single vacancy on a $B$-site of the kagome lattice (Fig. 4). For $J_{1}=0$, the presence or absence of this spin has strictly no effect on the ground state manifold, but switching on $J_{1}$ lifts the degeneracy. The $A$ and $C$ spins of the four first neighbors of the impurity experience an apparent magnetic field of strength $J_{1}$, in the direction opposite to the missing spin[35]. In a Néel antiferromagnet, parallel and perpendicular magnetic susceptibilities differ: $\mathcal{S}_{\parallel}=0$ while $\mathcal{S}_{\perp}>0$: the spins tend to orient perpendicularly to the magnetic field and to minimize their energy by a canting that decreases when moving away from the impurity. As a consequence, the $A$ and $C$ Néel ordered sub-lattices orient quasi-perpendicularly to the $B$ spins, with angles depending on the distance to the impurity (see correlations on Fig. 4) and alternating signs, similarly to Friedel oscillations already evidenced in other continuous spin models with single vacancies[36]. This phenomena is similar to the selection of the least-collinear order[7, 24], also observed on the $J_{1}J_{2}$HS model[20, 37, 38], explained by an effective positive biquadratic interaction $(\mathbf{S}_{i}\cdot\mathbf{S}_{j})^{2}$ between neighboring spins. However, with a unique impurity, the $A$ and $C$ quasi-Néel orders remain mutually independent and we still expect them to align collinearly at inifinitesimal temperatures, a partial survival of the initial ObTD. With one impurity, we have only two perpendicular spin directions, which is not enough to specify a chirality. On the $J_{1}J_{2}$HS, the only two independent Néel orders prevent the apparition of a $\mathbb{Z}_{2}$ chirality for Heisenberg spins: only a vectorial chirality can be defined, which is a continuous order parameter. However, with $XY$ spins, the chirality is scalar (along $z$) and gives rise to an Ising-like phase transition[6]. In the $J_{1}-J_{3}$HK model, with 3 sublattices and with vacancies on all types of sites, the 3 Néel orders tend to orient mutually perpendicularly, allowing to determine a local spin triedra with a scalar chirality.
@@
The phase diagram TpT-p is displayed in Fig. 1: the red dot at $p=0$ corresponds to the already known phase transition of the pure model, belonging to the universality class of the 4-state Potts model[26] (with critical exponents verifying $\alpha/\nu=1$, $\gamma/\nu=7/4$, and $1/\nu=3/2$), as clearly shown by finite-size scaling analysis (FSSA), while the blue dots is the emergent chiral transition with the expected exponents in the Ising universality class $\alpha/\nu=0$, $\gamma/\nu=7/4$, and $1/\nu=1$. We use the following relations for the maxima of the specific heat $C_{V}$ and of the (connected) susceptibility $\mathcal{S}_{X}$ of the order parameter $X$ (with $X=\Sigma$ or $\chi$) to determine the exponents of the phase transitions[41]:
@@
$\begin{split}C_{V}^{\rm max}(L)&=\frac{\overline{\langle\mathcal{H}^{2}\rangle}-\overline{\langle\mathcal{H}\rangle^{2}}}{NT^{2}}\simeq aL^{\alpha/\nu}+b,\\ \mathcal{S}_{X}^{\rm max}(L)&=\frac{\overline{\langle X^{2}\rangle}-\overline{\langle X\rangle^{2}}}{NT}\simeq cL^{\gamma/\nu}+d,\\ T^{Y^{\rm max}}(L)&\simeq eL^{-1/\nu}+T_{c},\end{split}$ (3)
@@
where brackets denote thermal averages and overlines are averages over disorder realizations. $T^{Y^{\rm max}}$ is the temperature of the maximum of $Y$, where $Y$ is $C_{V}$, $\mathcal{S}_{\Sigma}$ or $\mathcal{S}_{\chi}$, and $T_{c}$ is the critical temperature.
@@
Figure 5: (left) Maxima of heat capacity $C_{V}^{\rm max}$ versus the linear lattice size $L$ and (right) $T_{c}^{C_{V}}(L)$ versus $L^{-3/2}$, for several dilution rates pp.
@@
Figure 6: (left) Order parameter $\Sigma$ (increasing and decreasing with $T$) and $\chi$ (immediately decreasing) versus $T$ for different dilutions and $L=96$. (right) Maximum of the susceptibilty $\mathcal{S}_{\chi}^{\rm max}$ versus $L^{1.75}$. The dotted lines correspond to linear fits.
We now discuss the phase diagram for $p\neq 0$. For low impurity rate $p<0.05$, we have at first sight all the indications of a phase transition connected to the $K_{4}$ transition at $p=0$ (green points on Fig. 1) in the same universality class as the pure model. $T^{C_{V}^{\rm max}}(L)$ obeys Eq. (3) with a $T_{c}$ that rapidly decreases with $p$ (Fig. 5 right) and we observe a linear dependence of $C_{V}^{\rm max}(L)$ (Fig. 5 left). However, this behavior of $C_{V}$ is only valid up to a $L^{*}$ that decreases with $p$ (compare $p=0.02$ and $p=0.03$). For $p\geq 0.04$, the amplitude $a$ cancels. As explained above, site dilution leads to the existence of local random fields that are expected to destroy the order of the pure model at the lower critical dimension d=2d=2[15]. According to the Imry-Ma argument, order is destroyed by the formation of domains, whose size increases exponentially when the impurity rate decreases, as $e^{C/W}$ where $C$ is a constant and $W$ the strength of the random field[14]. Thus we expect that the disappearence of the $K_{4}$ phase transition due to dilution can only be observed at large lattice sizes. Note that other studies also report signs of transitions in the presence of effective random fields. On the $J_{1}J_{2}$HS model, a transition similar to the $p=0$ one is observed in presence of site-disorder [38], but not in Kunwar et al[19] where the authors do not go below $p=1/8$. However, by modifying the nature of the quenched disorder they observe an Imry-Ma contradicting transition[19] (see footnote 30), arguing that their lattice-symmetry breaking effect is very small[19]. A final example occurs, but in 3d model[25], where an effective gap at $T>0$ is invoked to overcome the Imry-Ma argument.
@@
At lower temperatures, a phase transition appears, where we observe simultaneously a rapid decrease of $\Sigma$ ($K_{4}$ order) and a rapid increase of $\chi$ (chiral order) as $T$ decreases (see Fig. 6 left). In the presence of site disorder, non-collinear configurations are energetically favored, overcoming the entropic advantage of the $K_{4}$ order at very low temperatures. The thermodynamic signature of this transition is the appearance of a small secondary maximum in $C_{V}$ at $T_{c}\simeq 4p$ for $p<0.05$, which increases weakly with $L$ (e.g., $C_{V}^{\rm max}(28)=1.15$, $C_{V}^{\rm max}(128)=1.20$). This prevents the determination of a reliable scaling exponent with FSSA, that however remains compatible with the scaling of the dilute Ising model[17, 42]. To obtain stable results in this regime, a disorder average over $20$ realizations is required. Results are also also limited to $L\leq 80$, due to thermalization difficulties.